On a school test, The average score for the girls in the class is 91, and the average score for the boys in the class is 85. If the average score for the class is 89, what fraction of the class are boys?
I should resolve it by a system...
91x + 85y= 89 (x+y)
But then??
Thanks in advance.

Answer by edjones(8007)

(91x+85y)/(x+y)=89
91x + 85y= 89 (x+y)
...=89x+89y
2x=4y
y=x/2
1/3 of the class are boys.
.
Ed